previous
next
Photos 6, page 28
next
previous
Lava Hike
April 23, 2010
Here's a view of the hike with an overlay of the map from the HVO Status Page.
And here are excerpts from the same page since my
last hike on April 8:
Photos aren't mine, they are from the
HVO image page.

Thursday, April 8, 2010 7:56 AM HST (Thursday, April 8, 2010 17:56 UTC)
Lava flowed through tubes and fed surface flows advancing slowly to the southeast down the
Pulama pali, east of the Royal Gardens
subdivision, as indicated by field observations and GOES satellite imagery. If the flows continue on their current path, they are projected to
rejoin the east edge of the TEB flow field west of the current viewing area.
Tuesday, April 20, 2010 8:26 AM HST (Tuesday, April 20, 2010 18:26 UTC)
Lava from the east rift zone vent, flowing into a lava tube system, continued to supply surface flows that have advanced slowly to the southeast
down the
Pulama pali, east of the Royal Gardens subdivision. HVO geologists in the field yesterday, mapping the flow margin, reported that
the flows continue their slow advancement on the coastal plain. The flow front was heading southeast along the east margin of the TEB flow
field, just west of the end of the
Kalapana access road, and was 415 yards (380 meters) northwest of the County viewing area. HVO scientists
will visit the flow field today to assess the current location of the flows.

Wednesday, April 21, 2010 8:36 AM HST (Wednesday, April 21, 2010 18:36 UTC)
Lava from the east rift zone vent, flowing into a lava tube system, continued to supply surface flows that have advanced slowly to the southeast
down the
Pulama pali, east of the Royal Gardens subdivision. HVO geologists in the field yesterday, mapping the flow margin, reported that
the flows continue their slow advancement on the coastal plain. The flow front was still heading southeast along the east margin of the TEB
flow field, just west of the end of the
Kalapana access road, and was 360 yards (330 meters) northwest of the County viewing area. The lava
was pushing through a thickly vegetated area (
kipuka), triggering small brush fires and minor methane bursts. HVO scientists will visit the flow
field today to assess the current location of the flows.
Monday, April 12, 2010 7:41 AM HST (Monday, April 12, 2010 17:41 UTC)
Lava from the east rift zone vent, flowing into a lava tube system, continues to supply surface flows that are advancing slowly to the southeast
down the
Pulama pali, east of the Royal Gardens subdivision. Field observations on Sunday afternoon indicated that the flows had reached
the 450 ft elevation, and are now about 1 mile from the east edge of the TEB flow field and a little over 2 miles from the coast. If the flows
continue on their current path, they are projected to rejoin the east edge of the TEB flow field west of the current viewing area.

Tuesday, April 13, 2010 7:41 AM HST (Tuesday, April 13, 2010 17:41 UTC)
Lava from the east rift zone vent, flowing into a lava tube system, continued to supply surface flows that are advancing slowly to the southeast
down the
Pulama pali, east of the Royal Gardens subdivision. GOES satellite imagery indicates that lava flows are approaching the coastal
plain, and within the next few days are likely to intersect the east edge of the TEB flow field west of the current viewing area.

Wednesday, April 14, 2010 7:37 AM HST (Wednesday, April 14, 2010 17:37 UTC)
Lava from the east rift zone vent, flowing into a lava tube system, continued to supply surface flows that are advancing slowly to the southeast
down the
Pulama pali, east of the Royal Gardens subdivision. Field observations made Tuesday afternoon indicate that the flows had
advanced 500 meters (1650 feet) since Sunday. Within the next few days, active lava flows are likely to intersect the east edge of the TEB flow
field west of the current viewing area.
Friday, April 16, 2010 7:38 AM HST (Friday, April 16, 2010 17:38 UTC)
Lava from the east rift zone vent, flowing into a lava tube system, continued to supply surface flows that are advancing slowly to the southeast
down the
Pulama pali, east of the Royal Gardens subdivision. Field observations on Thursday indicated that the active flows were within 300
meters (1000 feet) of rejoining the east margin of the TEB flow field on the coastal plain. At the current rate of advance, the lava is expected to
reach the coastal plain on Friday afternoon or Saturday morning.

Saturday, April 17, 2010 7:07 AM HST (Saturday, April 17, 2010 17:07 UTC)
Lava from the east rift zone vent, flowing into a lava tube system, continued to supply surface flows that have advanced slowly to the southeast
down the
Pulama pali, east of the Royal Gardens subdivision. The flows were very close to rejoininig the older TEB flow field on the coastal
plain on Thursday, and may have done so by now. HVO scientists will visit the flow field later today to assess the current location of the flows.

Sunday, April 18, 2010 9:23 AM HST (Sunday, April 18, 2010 19:23 UTC)
Lava from the east rift zone vent, flowing into a lava tube system, continued to supply surface flows that have advanced slowly to the southeast
down the
Pulama pali, east of the Royal Gardens subdivision. HVO geologists in the field yesterday, mapping the flow margin, reported that
the flow front had reached the east margin of the older TEB flow field on the coastal plain, about 0.5 miles (750 meters) northwest of the
County viewing area. HVO scientists will visit the flow field today to assess the current location of the flows. The ongoing deflation may reduce
the lava supply and advancement rate over the next day.

Monday, April 19, 2010 8:23 AM HST (Monday, April 19, 2010 18:23 UTC)
[The date of the overlay above]
Lava from the east rift zone vent, flowing into a lava tube system, continued to supply surface flows that have advanced slowly to the southeast
down the
Pulama pali, east of the Royal Gardens subdivision. HVO geologists in the field yesterday, mapping the flow margin, reported that
the flows continue their slow advancement on the coastal plain. The flow front had reached the vegetated area (
kipuka) just west of the end of
the
Kalapana access road, and were 0.3 miles (500 meters) northwest of the County viewing area. HVO scientists will visit the flow field today
to assess the current location of the flows.
15 April 2010
Visible/thermal image showing active flows on the pali
Flows continue to be active on the pali and are advancing southward. The flow front this morning was
approximately 1.4 km (0.9 miles) north of the County viewing area. Letters
A and B denote
corresponding points in the photograph (top) and thermal image (bottom). The active flows show up
clearly in the thermal image.
20 April 2010
Left.
Surface flows on the coastal plain were
active a couple hundred meters (yards) from
the current viewing trail at the end of Highway
130. The flows are burning along the margin
of a large
kipuka mauka of the viewing trail.

Right. While much of the flow stayed along
the margin of the
kipuka, a few lobes
ventured into the vegetation.
continued